OpenStudy (anonymous):

How to integrate this? ∫(8(x^.5)/(x^2+1)^.5)dx

OpenStudy (anonymous):

have you tried direct substitution?

OpenStudy (anonymous):

do you mean u-substitution?

OpenStudy (anonymous):

yes, the degree of the numerator is one less then that of the denominator, indicating some sort of ln relationship

OpenStudy (anonymous):

not sure what to sub...

OpenStudy (anonymous):

well, it may have been worth a try...i reach a dead end that way. Are you learning trig substitutions?

OpenStudy (anonymous):

x^2+1=u 2x=u' then...

OpenStudy (anonymous):

yeah i was thinking trig substitution, but it has a root x term at the top and i'm not sure how to deal with that using trig sub.

OpenStudy (anonymous):

Try: \[x = \tan \theta\] This comes from suggested substitutions when you see particular terms under a root sign, leading to pythagorean simplification. first table on first page http://www.stewartcalculus.com/data/CALCULUS%20Concepts%20and%20Contexts/upfiles/3c3-TrigonometSubstitu_Stu.pdf

OpenStudy (anonymous):

not sure how to apply it for this case since it has a root at the top and bottom...

OpenStudy (anonymous):

This is messed. sure the questions is right? this integral doesn't even want to exist in any tables of integrals i can find

OpenStudy (anonymous):

yeah, i'm pretty sure The original question was to find the volume of the solid when you rotate the area between the two curves: y= \[\frac{ 4 }{ x^{2}+1 }\] and y=\[\sqrt{x}\] for the region 0<x<2.3 This question came from expanding the integral

OpenStudy (anonymous):

rotate about the x axis i mean

OpenStudy (anonymous):

so, the the upper functions change in the middle of that range. You're doing disc?

OpenStudy (anonymous):

they intersect at 2.3ish, and i'm using disks

OpenStudy (anonymous):

They intersect at appox 1.504 https://www.desmos.com/calculator/qssiomirve

OpenStudy (anonymous):

ohhh, i left a square root out, my bad. the denominator of the first function is suppose to have a square root!

OpenStudy (anonymous):

\[\frac{ 4 }{ \sqrt{x ^{2}+1} }\]

OpenStudy (anonymous):

you may have used the shell equation for your integral, by using the disc, which is pi*r^2, the sqrt signs are destroyed